You just bought a 7-year zero coupon bond for $742.77. What is the (expected) taxable capital gain on this bond a year from now?

Question 3 options:

$14.33

$18.51

$22.47

$32.23

None of the above.
Answer #1

Given that,

price of a 7 year coupon bond = $742.77

So, yield on this bond = (FV/Price)^(1/t) - 1 = (1000/742.77)^(1/7) - 1 = 4.34%

So, price of the bond next year = Price now*(1+yield) = 742.77*(1+0.0434) = $775.00

So, capital gain = price next year - current price = 775 - 742.77 = $32.23

So, taxable capital gain - $32.23

Option D is correct.